OpenStudy (anonymous):

What is the probability of drawing yellow marble followed by a red marble from a bag containing 12 yellow marbles, 14 red marbles, and 15 green marbles if the first marble is not replaced?

OpenStudy (anonymous):

Total marbles = 12 + 14 + 15 = 41 Probability of drawing a yellow marble: 12/41 Probability of drawing a red marble without replacing the first marble: 14/40 Total probability = (12/41)*(14/40)